OpenStudy (anonymous):

A FM signal Vfm= 400sin[ 4pi x (10^10)t + 2 sin(3pi x 10^5t), is applied to a 1kilo ohm antenna what is the carrier frequency and the intelligence frequency? what is the modulation index? Determine the bandwidth using two methods what is the total power? power of the carrier? power of each sidebands?

OpenStudy (anonymous):

wc=4pi*10^10, so carrier frequency=2*10^10; and intelligence frequency=1.5*10^10;

OpenStudy (anonymous):

What is the formula for this. Please explain.